SPECTROSCOPY AND PREDISSOCIATION DYNAMICS OF THE Li_{2} F^{1}Σ_{g}^{+} STATE

\underline{S. ANTONOVA}, K. URBANSKI, G. LAZAROV, A. M. LYYRA, Department of Physics, Temple University, Philadelphia, PA 19122; L. LI, Department of Modern Applied Physics, Tsinghua University, Beijing, China.